11-14 Highland Automotive wishes to forecast the number

of new cars that will be sold next week. The following

table summarizes the number of new cars sold

during each of the past 12 weeks: (PLEASE ANSWER BY USING EXCEL)

week number sold
1 22
2 26
3 23
4 27
5 21
6 25
7 28
8 26
9 29
10 29
11 27
12 31

(a) Provide a forecast by using a 3-week weighted

moving average technique with weights 5, 3, and

1 (5 = most recent).

(b) Forecast sales by using an exponential smoothing

model with a = 0.45.

(c) Highland would like to forecast sales by using

linear trend analysis. What is the linear equation

that best fits the data?

(d) Which of the methods analyzed here would you

use? Explain your answer.
Answer #1

a)

A B E Week actuals wt moving avg 0.111111111 1 0.555555556 0.333333333 22 Wt Wt-1 Wt-2 2 1 2 26 3 23 Error МАРЕ 4 3 Sq Err 23

Formula under each cell

A E F G 3/9 -1/9 wt moving avg -5/9 1 Week actuals 2 1 22 Wt-2 Wt Wt-1 26 3 2 23 MAPE 4 3 Error Sq Err =ABS(E5) / A5 27 -SE$1

b)

A D G J M 24 |alpha 25 0.45 0.45 0.45 0.45 26 Week Sales SqError Error МАРЕ 22 27 26 2 28 22 -4 16 0.153846154 3 23 29 22 -1

Formula of the above excel calc

D G M A 26 Week Sales Error SqError МАРЕ 22 27 1 =D28- $B28 -ABS(G28)/$B28 28 2 26 =B27 -G28^2 D29-$B29 -ABS(G29)/SB29 23 =SD

D G M A 26 Week Sales Error SqError МАРЕ 22 27 1 =D28- $B28 -ABS(G28)/$B28 28 2 26 =B27 -G28^2 D29-$B29 -ABS(G29)/SB29 23 =SD

c) Performed the regression analysis below for the linear trend in excel's data analysis tool pack

A C D F F G H 1 SUMMARY OUTPUT 2 3 Regression Statistics Multiple 0.745499 R Square 6 Adjusted 0.511346 4 5 0.555769 7 Standa

d)If we compare a b and c the least sum of squares of error is in the case of exponential smoothening.

Hence we will use exponential smoothening for forecasting

Hope the above answer has helped you in understanding the problem. Please upvote the ans if it has really helped you. Good Luck!!